OpenStudy (anonymous):

SEVEN MORE THAN THREE TIMES A NUMBER IS 16

OpenStudy (anonymous):

3n+7=16 n=3

OpenStudy (sleepyjess):

3x+7=16 First you will need to subtract 7 from both sides 3x=9 Now divide each side by 3 x=?

OpenStudy (anonymous):

seven more then the quotinet of a number and -2 is 6

OpenStudy (anonymous):

3x+7=16 3x=9 x=3

OpenStudy (anonymous):

(n/-2)+7 n = 1/2

OpenStudy (anonymous):

thanks booboos

OpenStudy (anonymous):

i can do these in 20 secs just give me all lol

OpenStudy (anonymous):

lol kayy hold on

OpenStudy (anonymous):

6 more then twice a mumber is 20